ons 10. Rationalize each denominator. Simplify. 5 2 – 73 b) na os 75EVE a) V3 + V13 V3 - V13
Answer #1

() , : S(24) (9) = Els tres 4-3 - = |10+5v5) 9 한 응을 + 2 8+ 8-20 EING - g) 3) - - = 지는 6 -8) h9 - 9 - SS - = ㅉ - 5052| 인5 1 SSe) : -7. SS + 22 T VS - 2E ?fzyz S.zi : .5(JS + act) (s) - zxvil) : 55.4.2 S-8 : XS - 30 -3 g) 3 B ,3 ,3.% | 5- 5-6 3 + E : (